# Fast LiDAR Upsampling using Conditional Diffusion Models

### v1

@@ Abstract
The search for refining 3D LiDAR data has attracted growing interest motivated by recent techniques such as supervised learning or generative model-based methods. Existing approaches have shown the possibilities for using diffusion models to generate refined LiDAR data with high fidelity, although the performance and speed of such methods have been limited. These limitations make it difficult to execute in real-time, causing the approaches to struggle in real-world tasks such as autonomous navigation and human-robot interaction. In this work, we introduce a novel approach based on conditional diffusion models for fast and high-quality sparse-to-dense upsampling of 3D scene point clouds through an image representation. Our method employs denoising diffusion probabilistic models trained with conditional inpainting masks, which have been shown to give high performance on image completion tasks. We introduce a series of experiments, including multiple datasets, sampling steps, and conditional masks, to determine the ideal configuration, striking a balance between performance and inference speed. This paper illustrates that our method outperforms the baselines in sampling speed and quality on upsampling tasks using the KITTI-360 dataset. Furthermore, we illustrate the generalization ability of our approach by simultaneously training on real-world and synthetic datasets, introducing variance in quality and environments.
@@ III-B Image-based inpainting
In this work, we build an upsampling pipeline for 3D LiDAR, by integrating the DDPM architecture of LiDAR data generation [4] and the inpainting framework inspired by Palette [7]. In the standard DDPMs, the data generation is learned as an iterative denoising process from latent variables $z$ (Gaussian noise) into data $x$, called reverse diffusion. Each denoising step is performed by a neural network such as U-Net. In contrast, our approach uses conditional observable pixels $\hat{x}$ that is combined into the input $z_{t}$ of the neural network for each reverse diffusion step $t$, as depicted in Fig. 2. The observable pixels $\hat{x}$ are blended into the latent variables $z_{t}$ according to the mask $m$ representing known and unknown region: $m\odot\hat{x}+(1-m)\odot z_{t}$. Although the upsampling technique [10] used in R2DM [4] also uses masked sampling, extra harmonization steps are required due to the objective gap between the training and sampling. For the loss function, we modify the one used in the R2DM [4], by making the loss only be calculated on the predicted masked area and not the entire image.
@@ III-D Training method
Similar to Nakashima and Kurazume [4] we built an upsampling pipeline that takes in a LiDAR point cloud projected to an image and upsamples the input to a $64\times 1024$ image. In addition, we added 3 other inpainting masks for multitask-learning as suggested by Saharia et al.  [7]. Saharia et al. showed that training Palette on several tasks made the model generalize better and perform as well if not better on several tasks.
@@ IV-A Settings
In addition to the R2DM baseline, we also compare our methods against well-established interpolation techniques (Nearest-neighbor, Bilinear, and Bicubic). We also compare our results against the LiDAR diffusion model based LiDARGen [3] on 4x upsampling. Lastly, we compare with simulation-based supervised completion methods: Simulation base LiDAR super-resolution (LiDAR-SR) [15], Implicit LiDAR Network (ILN) [13], and Local Implicit Image Function (LIIF) [16].
@@ IV-D Results
Training task
@@
Ours A 8 KITTI-360 2-8x 0.6578 3.4690 0.0407 0.0858 44.27
Ours B 8 KITTI-360 2x 8.7424 15.9029 0.1778 0.2652 3.81
Ours C 8 KITTI-360 4x 0.6431 3.3912 0.0402 0.0844 45.55
Ours D 8 KITTI-360 8x 1.0010 3.9631 0.0554 0.1103 30.58
Ours E 8 KITTI-360 4x 0.8016 3.6900 0.0460 0.0934 37.50
Ours F 8 KITTI-360 2-8x 0.6868 3.4619 0.0422 0.0880 42.44
Ours G 8 KITTI-360 4x 0.6845 3.4757 0.0420 0.0876 42.66
Ours H 8 KITTI-360 4x 0.7472 3.6017 0.0433 0.0900 40.33
Ours I 8 KITTI-360 4x 0.7238 3.5826 0.0433 0.0896 39.64
Ours J 8 All datasets 4x 0.6299 3.3474 0.0422 0.0867 44.58
@@
  • To keep the table readable we have abbreviated each mask combination. Each letter is related to each mask: Simple (S), Jitter (J), Pepper noise (P), and Upsampling (U). Under the Upsampling column, we have included the upsampling ratios.

### v2

@@ Abstract
The search for refining 3D LiDAR data has attracted growing interest motivated by recent techniques such as supervised learning or generative model-based methods. Existing approaches have shown the possibilities for using diffusion models to generate refined LiDAR data with high fidelity, although the performance and speed of such methods have been limited. These limitations make it difficult to execute in real-time, causing the approaches to struggle in real-world tasks such as autonomous navigation and human-robot interaction. In this work, we introduce a novel approach based on conditional diffusion models for fast and high-quality sparse-to-dense upsampling of 3D scene point clouds through an image representation. Our method employs denoising diffusion probabilistic models trained with conditional inpainting masks, which have been shown to give high performance on image completion tasks. We introduce a series of experiments, including multiple datasets, sampling steps, and conditional masks. This paper illustrates that our method outperforms the baselines in sampling speed and quality on upsampling tasks using the KITTI-360 dataset. Furthermore, we illustrate the generalization ability of our approach by simultaneously training on real-world and synthetic datasets, introducing variance in quality and environments.
@@ III-B Image-based inpainting
In this work, we build an upsampling pipeline for 3D LiDAR, by integrating the DDPM architecture of LiDAR data generation [4] and the inpainting framework inspired by Palette [7]. In the standard DDPMs, the data generation is learned as an iterative denoising process from latent variables $z$ (Gaussian noise) into data $x$, called reverse diffusion. Each denoising step is performed by a neural network such as U-Net. In contrast, our approach combines conditional observable pixels $\hat{x}$ into the input $z_{t}$ of the neural network for each reverse diffusion step $t$, as depicted in Fig. 2. The observable pixels $\hat{x}$ are blended into the latent variables $z_{t}$ according to the mask $m$ representing known and unknown region: $m\odot\hat{x}+(1-m)\odot z_{t}$. Although the upsampling technique [10] used in R2DM [4] also uses masked sampling, extra harmonization steps are required due to the objective gap between the training and sampling. For the loss function, we modify the one used in the R2DM [4], by making the loss only be calculated on the predicted masked area and not the entire image.
@@ III-D Training method
Similar to Nakashima and Kurazume [4] we built an upsampling pipeline that takes in a LiDAR point cloud projected to an image and upsamples the input to a $64\times 1024$ image. In addition, we added 4 inpainting masks for multitask-learning as suggested by Saharia et al.  [7]. Saharia et al. showed that training Palette on several tasks made the model generalize better and perform as well if not better on several tasks.
@@ IV-A Settings
In addition to the R2DM baseline, we also compare our methods against well-established interpolation techniques (Nearest-neighbor, Bilinear, and Bicubic). We also compare our results against the LiDAR diffusion model based LiDARGen [3] on 4x upsampling. Lastly, we compare with simulation-based supervised completion methods: Simulation base LiDAR super-resolution (LiDAR-SR) [15], and Implicit LiDAR Network (ILN) [13].
@@ IV-B Evaluation metrics
As a considerable section of each sample keeps the original data, each metric is only calculated for the generated area decided by the mask to measure improvements effectively.
@@ IV-D Results
Training task Depth Reflectance
@@
Ours A 8 KITTI-360 $\times^{\ddagger}$ 0.6578 3.4690 0.0407 0.0858 44.27
Ours B 8 KITTI-360 2$\times$ 8.7424 15.9029 0.1778 0.2652 3.81
Ours C 8 KITTI-360 4$\times$ 0.6431 3.3912 0.0402 0.0844 45.55
Ours D 8 KITTI-360 8$\times$ 1.0010 3.9631 0.0554 0.1103 30.58
Ours E 8 KITTI-360 4$\times$ 0.8016 3.6900 0.0460 0.0934 37.50
Ours F 8 KITTI-360 2-8$\times$ 0.6868 3.4619 0.0422 0.0880 42.44
Ours G 8 KITTI-360 4$\times$ 0.6845 3.4757 0.0420 0.0876 42.66
Ours H 8 KITTI-360 4$\times$ 0.7472 3.6017 0.0433 0.0900 40.33
Ours I 8 KITTI-360 4$\times$ 0.7238 3.5826 0.0433 0.0896 39.64
Ours J 8 All datasets 4$\times$ 0.6299 3.3474 0.0422 0.0867 44.58
@@
  • †To keep the table readable we have abbreviated each mask combination. Each letter is related to each mask: Simple (S), Jitter (J), Pepper noise (P), and Upsampling (U). Under the Upsampling column, we have included the upsampling ratios.
  • ‡2-8x states that samples during training were randomly masked with one of the optional 2x, 4x, and 8x upsampling rates respectively.